2b) Prove or disprove the following statement. (2)

“If m divides a

n −b

n

, then m divides abn −ban

also.”


1
Expert's answer
2021-04-13T16:21:54-0400

Solution:

Given that "m" divides "a^n-b^n" .

We know that for any integer "n",

"a^n-b^n=(a-b)(a^{n-1}+a^{n-2}b+...+ab^{n-2}+b^{n-1})" ...(i)

"\Rightarrow m" divides "(a-b)".

Now, consider "ab^n-ba^n"

"=-ab(a^{n-1}-b^{n-1})"

"=-ab(a-b)(a^{n-2}+a^{n-3}b+...+ab^{n-3}+b^{n-2})" [Using (i)]

We can see that "(a-b)" is also a factor here and thus, divisible by "m".

Hence, yes, "m" divides "ab^n-ba^n".

Proved.